OpenStudy (anonymous):

WILL GIVE MEDAL!! NEED HELP ASAP!!! Simplify. 48 ÷ ((2 + 6) · 2) - 1

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

PEMDAS 48 ÷ ((2 + 6) · 2) - 1 Parenthesis first, add: 48 ÷ (8 · 2) - 1 Parenthesis, multiply: 48 ÷ 16 - 1 Division, divide: 3 - 1 Now subtract, what do you get? @birvin73

OpenStudy (igreen):

Subtract 3 - 1. @birvin73

OpenStudy (anonymous):

Parenthesis Exponents Multiplication Division Addition Subtraction 48/((2+6)*2)-1 2+6=8 48/(8*2)-1 8*2=16 48/16-1 48/16 =3 3-1?? what do u get?

OpenStudy (anonymous):

2

OpenStudy (igreen):

Yep, so your answer is 2.

OpenStudy (igreen):

Yep, that's how you should answer questions. @roz_mar23 Don't give them the answer, just guide them :)

OpenStudy (anonymous):

Thanx @iGreen

OpenStudy (anonymous):

igreen can you give roz_mar23 a medal

OpenStudy (igreen):

I already did :)

OpenStudy (igreen):

No problem. @roz_mar23

OpenStudy (anonymous):

can someone halp me